# The Automorphism Equivariant Hitchin Index

### v1

@@ 2.2 Hecke Transformations and Descent
Let $i\in\mathcal{I}_{0}$. We introduce the following notations. Denote by $M_{i}$ the set of isomorphism classes of $\langle f\rangle$-equivariant quasi parabolic rank two Higgs bundles $(E,F,\Phi)\rightarrow(X,\mathcal{D}_{i})$ with determinant $\langle f\rangle$-equivariantly isomorphic to $\Lambda,$ which satisfies the following conditions. For all $x\in X^{f}$, the eigenvalues of $f_{E}(x)$ are $\sqrt{\alpha_{x}}\zeta^{i_{x}}$ and $\sqrt{\alpha_{x}}\zeta^{-i_{x}}$, and $F_{x}\subset E_{x}$ is equal to the $\sqrt{\alpha_{x}}\zeta^{-i_{x}}$-eigenspace. Denote by $M_{i}^{\prime}$ the set of isomorphism classes of $\langle f\rangle$-equivariant quasi parabolic rank two Higgs bundles $(E,F,\Phi)\rightarrow(X,\mathcal{D}_{i})$ with determinant $\langle f\rangle$-equivariantly isomorphic to $\pi^{*}(\tilde{\Lambda}_{i})$ and such that $f_{E}$ acts as the identity on the fibres of $E$ above $X^{f}$. We will now show that the map given in equation (33) gives a bijection $\Psi_{i}:M_{i}\rightarrow M_{i}^{\prime}$. First of all, we note that the map $\Psi_{i}$ is well-defined since Hecke transformations at different points $x\in\mathcal{D}_{i}$ commutes. Let $(E_{0},F_{0},\Phi)\in M_{i}$. Define the flags $F_{0}(S_{i})=((F_{0})_{x}\otimes(S_{i})_{x})_{x\in X^{f}}$. Then $(E_{1},F_{1},\Phi):=(E_{0}(S_{i}),F_{0}(S_{i}),\Phi)$ canonically defines a $\langle f\rangle$-equivariant quasi parabolic Higgs bundle as in Proposition 2.4 with eigenvalues $\{1,\zeta^{-2i_{x}}\}_{x\in\mathcal{D}_{i}}$. For all $x\in\mathcal{D}_{i}$ set $b_{i}(x)=[2m_{x}i_{x}]_{p}$. Applying at each $x\in\mathcal{D}_{i}$ a total of $b_{i}(x)$ Hecke modifications to $(E_{1},F_{1},\Phi_{1})$, we obtain a new $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle $(E_{2},F_{2},\Phi_{2})=\prod_{x\in\mathcal{D}_{i}}\Psi_{x}^{b_{i}(x)}(E_{1},\Phi_{1})$. By construction, the $\langle f\rangle$-action is trivial on $E_{\mid X^{f}}$. This fact follows from Proposition 2.4. We now show $\det(E_{2})\simeq\pi^{*}(\Lambda_{i})$. As each Hecke transform at $x$ tensors the determinant by $(-x)$, we see that the determinant of $E_{2}$ is given by
$\begin{split}&\det(E_{2})=\det(E_{1})\left(-\sum_{x\in\mathcal{D}_{i}}b_{i}(x)\cdot x\right)=\Lambda\left(2S_{i}-\sum_{x\in\mathcal{D}_{i}}b_{i}(x)\cdot x\right)\\ &=\Lambda\left(2B+\sum_{x\in\mathcal{D}_{i}}(2[m_{x}i_{x}]_{p}-[2m_{x}i_{x}])\cdot x\right).\end{split}$ (34)
@@
Recall that $D_{i}=\{x\in\operatorname{\tilde{\mathcal{D}}}_{i},[m_{x}i_{x}]>(p-1)/2\}$. If we have $[m_{x}i_{x}]\leq(p-1)/2$ then $x\notin D_{i}$ and we have $2[m_{x}i_{x}]_{p}=[2m_{x}i_{x}]_{p}$. If not, then $\tilde{x}\in D_{i}$ and one can use that $p$ is odd to show that $[2m_{x}i_{x}]_{p}=2[m_{x}i_{x}]_{p}-p.$ Thus
$\begin{split}&\sum_{x\in\mathcal{D}_{i}}(2[m_{x}i_{x}]_{p}-[2m_{x}i_{x}])\cdot x=\sum_{x\in\mathcal{D}_{i}:[m_{x}i_{x}]>(p-1)/2}p\cdot x=\pi^{*}(D_{i}).\end{split}$ (35)
@@
Combining equation (34) with equation (35), and using $\pi^{*}(\tilde{\Lambda})=\Lambda(2B)$, we obtain the desired isomorphism $\det(E_{2})\simeq\pi^{*}(\tilde{\Lambda}_{i}).$ Thus it only remains to verify that $\Phi_{2}$ is holomorphic on $X\setminus\mathcal{D}_{i}$ with a pole of order at most one at each $x\in\mathcal{D}_{i}$ and with residue nilpotent with respect to the flag. For each tuple $a=(a_{x})_{x\in\mathcal{D}_{i}}$ with $0\leq a_{x}<b_{i}(x)$ define
@@
It follows from the implication (22), that the Higgs field of $\operatorname{\mathcal{E}}_{a}$ is in fact holomorphic on $X$. Thus, it is only after we apply the last product of Hecke modifications that the resulting Higgs field $\Phi_{2}$ aquires a potential pole divisor, which will be a subset of $\mathcal{D}_{i}$ by construction, and each pole will be of order at most one, with residue nilpotent with respect to the flag (this last fact was shown above and follows from (27)). Thus we have shown that $\Psi_{i}$ defines a bijection $M_{i}\rightarrow M_{i}^{\prime}$.
@@
Descent for holomorphic vector bundles is discussed in detail in [5]. We now extend descent to Higgs bundles. Denote by $\tilde{M}_{i}$ the set of isomorphism classes of quasi-parabolic rank two Higgs bundles on $(\tilde{X},\operatorname{\tilde{\mathcal{D}}}_{i})$ with fixed determinant $\tilde{\Lambda}_{i}$. We will show that pullback with respect to the quotient map $\pi:X\rightarrow\tilde{X}$ induces a bijection $\pi^{*}:\tilde{M}_{i}\rightarrow M_{i}^{\prime}$ with inverse given by descent $\operatorname{\mathcal{E}}\mapsto\operatorname{\mathcal{E}}/\langle f\rangle$. Define the $\langle f\rangle$-invariant divisor $R_{i}=\mathcal{D}_{i}+(1-p)(X^{f}\setminus\mathcal{D}_{i})$. First we observe that there is a canonical $\langle f\rangle$-equivariant isomorphism
@@
$\Phi^{\prime}/\langle f\rangle\in H^{0}(\tilde{X},\operatorname{End}_{0}(E^{\prime}/f)\otimes K_{\tilde{X}}(\operatorname{\tilde{\mathcal{D}}}_{i})).$
Thus the triple $(E^{\prime}/\langle f\rangle,F^{\prime}/\langle f\rangle,\Phi/\langle f\rangle)$ defines a quasi parabolic Higgs bundle on $\tilde{X}$, which we denote by $\operatorname{\mathcal{E}}^{\prime}/\langle f\rangle\in\tilde{M}_{i}.$ We now show (38). By the first part of the proof, we see that this is equivalent to proving that for all $\operatorname{\mathcal{E}}=(E,F,\Phi)\in M_{i}$, we have that
@@
By (19) we have equality between (41) and (42), and this gives
@@
Evaluating at $z=0$ and using that $f_{l,k}^{i,j}(0)=\delta_{l,k}^{i,j},$ (where the right hand side is Kronecker’s delta function) we obtain that $\phi_{i,j}(0)=0$ Assume inductively that we have shown that $\phi^{(m)}_{l,k}(0)=0$ for all $(l,k)$ and all $m=0,1,2,...,n-1$ with $n\leq p-2$. Differentiating equation (43) $n$ times and evaluating at $z=0$ and using that $\frac{\partial f}{\partial z}(z)=\zeta^{n_{x}}$, we get
@@
Substituting equation (49) into equation (47) and equation (49) into equation (48) and then computing the difference $\mu_{\operatorname{Par}}(\varXi_{i}(\operatorname{\mathcal{E}}))-\mu_{\operatorname{Par}}(\varXi_{i}(H))$ gives
@@
Thus we have proven equation (46). We now use this to prove $\operatorname{Im}(\varXi_{i})=\tilde{\mathcal{M}}_{i}$. Observe that for any $(E,F,\Phi)\in M_{i}^{\prime}$, there is an induced bijection between $\Phi$-invariant and $\langle f\rangle$-equivariant line subbundles $H$ of $E$ and $\Phi/\langle f\rangle$-invariant line subbundles of $E/\langle f\rangle$, and this is given by $H\mapsto H/\langle f\rangle$. This fact and equation (46) implies $\operatorname{Im}(\varXi_{i})\subset\tilde{\mathcal{M}}_{i}.$ Now let $\operatorname{\mathcal{E}}=(E,\Phi)\in M_{i}$ and assume $\varXi_{i}(\operatorname{\mathcal{E}})\in\tilde{\mathcal{M}}_{i}$. Assume towards a contradiction that $\operatorname{\mathcal{E}}$ is not stable. Since $\deg(\Lambda)$ is odd, this is equivalent to $\operatorname{\mathcal{E}}$ not being semi-stable. Consider the Higgs Narasimhan-Harder filtration $0\subsetneq H\subsetneq E$. Observe that $f_{E}(f^{*}(H))$ is a line subbundle which also induce a Higgs Narasimhan-Harder filtration, and therefore we must have $f_{E}(f^{*}(H))=H$ by uniqueness. Therefore $H$ is $\langle f\rangle$-equivariant and $\Phi$-invariant. Hence we can apply equation (46) to obtain a contradiction. Thus $\operatorname{\mathcal{E}}$ is stable. We now argue that $\varXi_{i}$ and its inverse are morphisms of varieties. Both the target and the source of $\varXi_{i}$ can be obtained as GIT quotients [57, 69, 70], and therefore it suffices to show that one can apply Hecke transformations in families. This was done for stable bundles in [5], and the analysis therein can readily be repeated for Higgs bundles. In particular, we can apply Hecke modications to $(\operatorname{\mathbbm{E}},\mathbb{\Phi})$ and we define
@@
It is a routine verification that $(\tilde{\operatorname{\mathbbm{E}}}_{i},\tilde{\mathbb{F}}_{i},\mathbb{\Phi}_{i})\rightarrow\tilde{\mathcal{M}}_{i}\times\tilde{X}$ is a $\mathbb{T}$-equivariant universal parabolic Higgs bundle, and that we have a canonical equivariant isomorphism as in equation (3). ∎
@@ 2.3 Identification of The Torus Fixed Locus
Let $M$ denote the set of odd integers $m$ with $31\leq m\leq 2g-$ 3. For each $m\in M$, let $\operatorname{\mathcal{F}}_{m}$ be the set of isomorphism classes of rank two Higgs bundles $(E,\Phi_{b})$ on $X$ of the following form.
@@
  • The Higgs field $\Phi_{b}$ is the off-diagonal Higgs field associated with a section $b\in H^{0}(X,L_{2}L_{1}^{-1}K_{X})$. The degree of the zero divisor of $b$ is equal to $m$.
For $m\in\mathbb{Z}_{\geq 0}$ let $X_{m}$ (resp. $\tilde{X}_{m}$) denote the space of positive effectice divisors of degree $m$ on $X$ (resp. $\tilde{X}$). Now fix $m\in M$. The following facts are shown by Hitchin in [40]. Every Higgs bundle in $\operatorname{\mathcal{F}}_{m}$ is stable, and thus $\operatorname{\mathcal{F}}_{m}$ is a naturally a subspace of $\mathcal{M}$. Each $\operatorname{\mathcal{F}}_{m}$ is connected and we have
$\mathcal{M}^{\mathbb{T}}=\mathcal{N}\sqcup_{m\in M}\operatorname{\mathcal{F}}_{m}.$ (57)
Moreover, the map $\delta:\operatorname{\mathcal{F}}_{m}\rightarrow X_{m}$, given by $(E,\Phi_{b})\mapsto(b)\in X_{m}$, gives $\operatorname{\mathcal{F}}_{m}$ the structure of a principal $\operatorname{Jac}(X)[2]$-bundle (with action given by tensoring). We notice that pullback with respect to $f$ naturally equips $\operatorname{\mathcal{F}}_{m},\ \operatorname{Jac}(X)[2]$ and $X_{m}$ with an $\langle f\rangle$-action.
Fix $m\in M$. Recall [39] that the universal Higgs bundle restricts to a direct sum of $\mathbb{T}$-equivariant line bundles $\mathbb{E}_{m,1}\oplus\mathbb{E}_{m,0}\rightarrow\operatorname{\mathcal{F}}_{m}\times X$, where $\mathbb{T}$ acts with weight $t$ on $\mathbb{E}_{m,1}$ and with weight equal to unity on $\mathbb{E}_{m,0}$. Fix $i\in\mathcal{I}$. We will now identity the components of $\mathcal{M}_{i}\cap\operatorname{\mathcal{F}}_{m}$ as well as the $\mathbb{T}\times\langle f\rangle$-weight subbundle decomposition of $\operatorname{\mathbbm{E}}$ restricted to these components.
@@
Set $\mathcal{J}_{m}=\sqcup_{i\in\mathcal{I}}\mathcal{J}_{i,m}$, where for each $i\in\mathcal{I}$, we define $\mathcal{J}_{i,m}$ to be the set of divisors of the form $j=\sum_{x\in X^{f}}j_{x}\cdot x,$ such that $l_{j}=(m-\deg(j))/p$ is a non-negative integer and for all $x\in X^{f}$, we have that $j_{x}\in\{0,...,p-1\}$ and
@@
Let $i\in\mathcal{I},m\in M$ and $j\in\mathcal{J}_{i,m}$. For all $x\in X^{f}$, it follows from equation (58) that there is a unique $\epsilon_{j,x}\in\{1,-1\}$ such that $\epsilon_{j,x}2i_{x}=-n_{x}(1+j_{x})\mod p.$
@@
The set of components of the $\langle f\rangle$-fixed locus $X_{m}^{f}$ is in bijection with $\mathcal{J}_{m}$, and the set of components of $\mathcal{M}_{i}\cap\mathcal{F}_{m}$ is in bijection with $\mathcal{J}_{i,m}.$ Fix $j\in\mathcal{J}_{i,m}$ and denote the associated components by $X^{f}_{m,j}$ and $\mathcal{M}^{\mathbb{T}}_{i,j}$ respectively.
@@
    We have an isomorphism $X^{f}_{m,j}\simeq\tilde{X}_{l_{j}}.$ Further, the morphism $\delta$ restricts to a projection of a principal $\operatorname{Jac}(\tilde{X})[2]$-bundle $\mathcal{M}^{\mathbb{T}}_{i,j}\rightarrow X^{f}_{m,j}$. Thus
    $\mathcal{M}_{i,j}^{\mathbb{T}}=\{(E,\Phi_{b})\in\mathcal{M}^{f}\cap\mathcal{F}_{m}:[\operatorname{ord}_{x}(b)]_{p}=j_{x},\ \forall x\in X^{f}\}.$ (59)
@@
    $\operatorname{\mathbbm{E}}_{\mid\mathcal{M}_{i,j}^{\mathbb{T}}\times\{x\}}=\operatorname{\mathbbm{E}}_{m,1}(x)\cdot\sqrt{\hat{\alpha}}_{x}\hat{\zeta}^{\epsilon_{j,x}i_{x}}\oplus\operatorname{\mathbbm{E}}_{m,0}(x)\cdot\sqrt{\hat{\alpha}}_{x}\hat{\zeta}^{-\epsilon_{j,x}i_{x}}.$ (60)

### v2

@@ 1.2 Hecke Transformations and Descent
Let $i\in\mathcal{I}_{0}$. We introduce the following notations. Denote by $M_{i}$ the set of isomorphism classes of $\langle f\rangle$-equivariant quasi parabolic rank two Higgs bundles $(E,F,\Phi)\rightarrow(X,\mathcal{D}_{i})$ with determinant $\langle f\rangle$-equivariantly isomorphic to $\Lambda,$ which satisfies the following conditions. For all $x\in X^{f}$, the eigenvalues of $f_{E}(x)$ are $\sqrt{\alpha_{x}}\zeta^{i_{x}}$ and $\sqrt{\alpha_{x}}\zeta^{-i_{x}}$, and $F_{x}\subset E_{x}$ is equal to the $\sqrt{\alpha_{x}}\zeta^{-i_{x}}$-eigenspace. Denote by $M_{i}^{\prime}$ the set of isomorphism classes of $\langle f\rangle$-equivariant quasi parabolic rank two Higgs bundles $(E,F,\Phi)\rightarrow(X,\mathcal{D}_{i})$ with determinant $\langle f\rangle$-equivariantly isomorphic to $\pi^{*}(\tilde{\Lambda}_{i})$ and such that $f_{E}$ acts as the identity on the fibres of $E$ above $X^{f}$.
We will now show that the map given in equation (14) gives a bijection $\Psi_{i}:M_{i}\rightarrow M_{i}^{\prime}$. First of all, we note that the map $\Psi_{i}$ is well-defined since Hecke transformations at different points $x\in\mathcal{D}_{i}$ commutes.
Let $(E_{0},F_{0},\Phi)\in M_{i}$. Define the flags $F_{0}(S_{i})=((F_{0})_{x}\otimes(S_{i})_{x})_{x\in X^{f}}$. Then $(E_{1},F_{1},\Phi):=(E_{0}(S_{i}),F_{0}(S_{i}),\Phi)$ canonically defines a $\langle f\rangle$-equivariant quasi parabolic Higgs bundle as in Proposition 1.4 with eigenvalues $\{1,\zeta^{-2i_{x}}\}_{x\in\mathcal{D}_{i}}$. Applying at each $x\in\mathcal{D}_{i}$ a total of $b_{i}(x)$ Hecke modifications to $(E_{1},F_{1},\Phi_{1})$, we obtain a new $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle $(E_{2},F_{2},\Phi_{2})=\prod_{x\in\mathcal{D}_{i}}\Psi_{x}^{b_{i}(x)}(E_{1},\Phi_{1})$. By construction, the $\langle f\rangle$-action is trivial on $E_{\mid X^{f}}$. This fact follows from Proposition 1.4.
We now show $\det(E_{2})\simeq\pi^{*}(\Lambda_{i})$. As each Hecke transform at $x$ tensors the determinant by (x)(-x), we see that the determinant of $E_{2}$ is given by
$\begin{split}&\det(E_{2})=\det(E_{1})\left(-\sum_{x\in\mathcal{D}_{i}}b_{i}(x)\cdot x\right)=\Lambda\left(2S_{i}-\sum_{x\in\mathcal{D}_{i}}b_{i}(x)\cdot x\right)\\ &=\Lambda\left(2B+\sum_{x\in\mathcal{D}_{i}}(2[m_{x}i_{x}]_{p}-[2m_{x}i_{x}]_{p})\cdot x\right).\end{split}$ (15)
@@
Recall that $D_{i}=\{x\in\operatorname{\tilde{\mathcal{D}}}_{i},[m_{x}i_{x}]_{p}>(p-1)/2\}$. If we have $m_{x}i_{x}]_{p}\leq(p-1)/2$ then $x\notin D_{i}$ and we have $2[m_{x}i_{x}]_{p}=[2m_{x}i_{x}]_{p}$. If not, then $\tilde{x}\in D_{i}$ and one can use that $p$ is odd to show that $2m_{x}i_{x}]_{p}=2[m_{x}i_{x}]_{p}-p.$ Thus
$\begin{split}&\sum_{x\in\mathcal{D}_{i}}(2[m_{x}i_{x}]_{p}-[2m_{x}i_{x}]_{p})\cdot x=\sum_{x\in\mathcal{D}_{i}:[m_{x}i_{x}]_{p}>(p-1)/2}p\cdot x=\pi^{*}(D_{i}).\end{split}$ (16)
Combining equation (LABEL:detE21) with equation (LABEL:eq:Dbeta), and using $\pi^{*}(\tilde{\Lambda})=\Lambda(2B)$, we obtain the desired isomorphism $\det(E_{2})\simeq\pi^{*}(\tilde{\Lambda}_{i}).$
@@
Thus it only remains to verify that $\Phi_{2}$ is holomorphic on $X\setminus\mathcal{D}_{i}$ with a pole of order at most one at each $x\in\mathcal{D}_{i}$ and with residue nilpotent with respect to the flag. For each tuple $a=(a_{x})_{x\in\mathcal{D}_{i}}$ with $0\leq a_{x}<b_{i}(x)$ define
@@
It follows from the implication (LABEL:implication), that the Higgs field of $\operatorname{\mathcal{E}}_{a}$ is in fact holomorphic on $X$. Thus, it is only after we apply the last product of Hecke modifications that the resulting Higgs field $\Phi_{2}$ aquires a potential pole divisor, which will be a subset of $\mathcal{D}_{i}$ by construction, and each pole will be of order at most one, with residue nilpotent with respect to the flag (this last fact was shown above and follows from (8)). Thus we have shown that $\Psi_{i}$ defines a bijection $M_{i}\rightarrow M_{i}^{\prime}$.
@@
Descent for holomorphic vector bundles is discussed in detail in [5]. We now extend descent to Higgs bundles. Denote by $\tilde{M}_{i}$ the set of isomorphism classes of quasi-parabolic rank two Higgs bundles on $(\tilde{X},\operatorname{\tilde{\mathcal{D}}}_{i})$ with fixed determinant $\tilde{\Lambda}_{i}$. We will show that pullback with respect to the quotient map $\pi:X\rightarrow\tilde{X}$ induces a bijection $\pi^{*}:\tilde{M}_{i}\rightarrow M_{i}^{\prime}$ with inverse given by descent $\operatorname{\mathcal{E}}\mapsto\operatorname{\mathcal{E}}/\langle f\rangle$.
Define the $\langle f\rangle$-invariant divisor $R_{i}=\mathcal{D}_{i}+(1-p)(X^{f}\setminus\mathcal{D}_{i})$. First we observe that there is a canonical $\langle f\rangle$-equivariant isomorphism
@@
$\Phi^{\prime}/\langle f\rangle\in H^{0}(\tilde{X},\operatorname{End}_{0}(E^{\prime}/\langle f\rangle)\otimes K_{\tilde{X}}(\operatorname{\tilde{\mathcal{D}}}_{i})).$
Thus the triple $(E^{\prime}/\langle f\rangle,F^{\prime}/\langle f\rangle,\Phi/\langle f\rangle)$ defines a quasi parabolic Higgs bundle on $\tilde{X}$, which we denote by $\operatorname{\mathcal{E}}^{\prime}/\langle f\rangle\in\tilde{M}_{i}.$
We now show (19). By the first part of the proof, we see that this is equivalent to proving that for all $\operatorname{\mathcal{E}}=(E,F,\Phi)\in M_{i}$, we have that
@@
By (2) we have equality between (22) and (LABEL:yh), and this gives
@@
Evaluating at $z=0$ and using that $f_{l,k}^{i,j}(0)=\delta_{l,k}^{i,j},$ (where the right hand side is Kronecker’s delta function) we obtain that $\phi_{i,j}(0)=0$ Assume inductively that we have shown that $\phi^{(m)}_{l,k}(0)=0$ for all $(l,k)$ and all m=0,1,2,,n1m=0,1,2,...,n-1 with $n\leq p-2$. Differentiating equation (24) $n$ times and evaluating at $z=0$ and using that $\frac{\partial f}{\partial z}(z)=\zeta^{n_{x}}$, we get
@@
Substituting equation (LABEL:qqq1) into equation (28) and equation (LABEL:qqq1) into equation (29) and then computing the difference $\mu_{\operatorname{Par}}(\varXi_{i}(\operatorname{\mathcal{E}}))-\mu_{\operatorname{Par}}(\varXi_{i}(H))$ gives
@@
[…]
We now argue that $\varXi_{i}$ and its inverse are morphisms of varieties. Both the target and the source of $\varXi_{i}$ can be obtained as GIT quotients [68, 80, 81], and therefore it suffices to show that one can apply Hecke transformations in families. This was done for stable bundles in [5], and the analysis therein can readily be repeated for Higgs bundles. In particular, we can apply Hecke modifications to $(\operatorname{\mathbbm{E}},\mathbb{\Phi})$ and we define
@@
It is a routine verification that $(\tilde{\operatorname{\mathbbm{E}}}_{i},\tilde{\mathbb{F}}_{i},\mathbb{\Phi}_{i})\rightarrow\tilde{\mathcal{M}}_{i}\times\tilde{X}$ is a $\mathbb{T}$-equivariant universal parabolic Higgs bundle, and that we have a canonical equivariant isomorphism as in equation (LABEL:mu_pequi). ∎ To see that (LABEL:uni) holds we argue as follows. For each $E\in\mathcal{M}_{i}$ and $x\in\mathcal{D}_{i}$ the flag $F_{x}\subset E_{x}$ as specified by Proposition 1.1 is given by the $\sqrt{\alpha_{x}}\zeta^{-i_{x}}$-eigenspace of the unique lift $f_{E}$ guaranteed by Proposition 1.1. It follows therefore from Remark 2 that $f_{\operatorname{\mathbbm{E}}}$ act with weight $\sqrt{\alpha_{x}}^{-1}\zeta^{i_{x}}$ on $\tilde{\mathbb{F}}_{i,x}$. As $\operatorname{\mathbbm{L}}_{i,x}=\operatorname{Hom}(\tilde{\mathbb{F}}_{i,x},(\tilde{\operatorname{\mathbbm{E}}}_{i,x}/\tilde{\mathbb{F}}_{i,x}))\cong\tilde{\mathbb{F}}_{i,x}^{*}\otimes(\tilde{\operatorname{\mathbbm{E}}}_{i,x}/\tilde{\mathbb{F}}_{i,x})$ this implies that (LABEL:uni) holds.
@@ 1.3 Identification of The Torus Fixed Locus
Let $C$ denote the set of odd integers $c$ with $1\leq c\leq 2g-3$. For each $c\in C$, let $\operatorname{\mathcal{F}}_{c}$ be the set of isomorphism classes of rank two Higgs bundles $(E,\Phi_{b})$ on $X$ of the following form.
@@
  • The Higgs field $\Phi_{b}$ is the off-diagonal Higgs field associated with a section $b\in H^{0}(X,L_{2}L_{1}^{-1}K_{X})$. The degree of the zero divisor of $b$ is equal to $c$.
For $c\in\mathbb{Z}_{\geq 0}$ let $X_{c}$ (resp. $\tilde{X}_{c}$) denote the space of positive effectice divisors of degree $c$ on $X$ (resp. $\tilde{X}$). Now fix $c\in C$. The following facts are shown by Hitchin in [47]. Every Higgs bundle in $\operatorname{\mathcal{F}}_{c}$ is stable, and thus $\operatorname{\mathcal{F}}_{c}$ is a naturally a subspace of $\mathcal{M}$. Each $\operatorname{\mathcal{F}}_{c}$ is connected and we have
$\mathcal{M}^{\mathbb{T}}=\mathcal{N}\sqcup_{c\in C}\operatorname{\mathcal{F}}_{c}.$ (38)
Moreover, the map $\delta:\operatorname{\mathcal{F}}_{c}\rightarrow X_{c}$, given by $(E,\Phi_{b})\mapsto(b)\in X_{c}$, gives $\operatorname{\mathcal{F}}_{c}$ the structure of a principal $\operatorname{Jac}(X)[2]$-bundle (with action given by tensoring). We notice that pullback with respect to $f$ naturally equips $\operatorname{\mathcal{F}}_{c},\ \operatorname{Jac}(X)[2]$ and $X_{c}$ with an $\langle f\rangle$-action.
Fix $c\in C$. Recall [45, proof of Lemma 6.1] that the universal Higgs bundle restricts to a direct sum of $\mathbb{T}$-equivariant line bundles $\mathbb{E}_{c,1}\oplus\mathbb{E}_{c,0}\rightarrow\operatorname{\mathcal{F}}_{c}\times X$, where $\mathbb{T}$ acts with weight $t$ on $\mathbb{E}_{c,1}$ and with weight equal to unity on $\mathbb{E}_{c,0}$. Fix $i\in\mathcal{I}$. We will now identity the components of $\mathcal{M}_{i}\cap\operatorname{\mathcal{F}}_{c}$ as well as the $\mathbb{T}\times\langle f\rangle$-weight subbundle decomposition of $\operatorname{\mathbbm{E}}$ restricted to these components.
@@
Set $\mathcal{J}_{c}=\sqcup_{i\in\mathcal{I}}\mathcal{J}_{i,c}$, where for each $i\in\mathcal{I}$, we define $\mathcal{J}_{i,c}$ to be the set of divisors of the form $j=\sum_{x\in X^{f}}j_{x}\cdot x,$ such that $l_{j}=(c-\deg(j))/p$ is a non-negative integer and for all $x\in X^{f}$, we have that $j_{x}\in\{0,...,p-1\}$ and
@@
Let $i\in\mathcal{I},c\in C$ and $j\in\mathcal{J}_{i,c}$. For all $x\in X^{f}$, it follows from equation (39) that there is a unique $\epsilon_{j,x}\in\{1,-1\}$ such that
$\epsilon_{j,x}2i_{x}=-n_{x}(1+j_{x})\mod p.$ (40)
@@
The set of components of the $\langle f\rangle$-fixed locus $X_{c}^{f}$ is in bijection with $\mathcal{J}_{c}$, and the set of components of $\mathcal{M}_{i}\cap\operatorname{\mathcal{F}}_{c}$ is in bijection with $\mathcal{J}_{i,c}.$ Fix $j\in\mathcal{J}_{i,c}$ and denote the associated components by $X^{f}_{c,j}$ and $\mathcal{M}^{\mathbb{T}}_{i,j}$ respectively.
@@
    We have an isomorphism $X^{f}_{c,j}\simeq\tilde{X}_{l_{j}}.$ Further, the morphism $\delta$ restricts to a projection of a principal $\operatorname{Jac}(\tilde{X})[2]$-bundle $\mathcal{M}^{\mathbb{T}}_{i,j}\rightarrow X^{f}_{c,j}$. Thus
    $\mathcal{M}_{i,j}^{\mathbb{T}}=\{(E,\Phi_{b})\in\mathcal{M}^{f}\cap\operatorname{\mathcal{F}}_{c}:[\operatorname{ord}_{x}(b)]_{p}=j_{x},\ \forall x\in X^{f}\}.$ (41)
@@
    $\operatorname{\mathbbm{E}}_{\mid\mathcal{M}_{i,j}^{\mathbb{T}}\times\{x\}}=\operatorname{\mathbbm{E}}_{c,1}(x)\otimes\sqrt{\hat{\alpha}_{x}}^{-1}\hat{\zeta}^{-\epsilon_{j,x}i_{x}}\oplus\operatorname{\mathbbm{E}}_{c,0}(x)\otimes\sqrt{\hat{\alpha}_{x}}^{-1}\hat{\zeta}^{\epsilon_{j,x}i_{x}}.$ (42)